A balanced-Y load is in parallel with a balanced-D-connected capacitor bank.

The Y load has an impedance ZY ¼ ð3 þ j4Þ W per phase, and its neutral is

grounded through an inductive reactance Xn ¼ 2 W. The capacitor bank has

a reactance Xc ¼ 30 W per phase. Draw the sequence networks for this load

and calculate the load-sequence impedances.


1
Expert's answer
2021-02-10T01:09:03-0500

"Z^1 = Zr || - Jxc"

"= \frac{(3+j4)-j\frac{30}{3}}{(3+j4)-j\frac{30}{3}}"

= 6.666 + j 3.33